OpenStudy (anonymous):

Which proportion would you use to find 38% of 90?

OpenStudy (anonymous):

38%= 38/100

OpenStudy (anonymous):

38%= 38/100

OpenStudy (anonymous):

ok say you were to find 40% of 70 40%= 40/100 so that would have to equal x/70 where x is your answer

OpenStudy (anonymous):

\[\frac{38}{100}=\frac{x}{90} \]